OpenStudy (anonymous):

my assignment has same diagrams and similar questions but the questions are different my assignment is speical angle types there are 3 assignments of speical angle types thats why i need help i know ive been asking alot cus i have ADHD so its hard for me to understand unlike other people so my question is what is measure angle 1 =

OpenStudy (anonymous):

attachment?

OpenStudy (anonymous):

First off, your instructor made a poor choice in calling it angle 1; it could confuse the reader into thinking that the angle is 1 degree.

OpenStudy (anonymous):

ik but that what is on the assignment its up there 7 times same equation measure angle 1 but differetn diagrams

OpenStudy (amistre64):

360 -2(125) ----------- = 180-125 = ang1; if i see it right 2

OpenStudy (amistre64):

180-(arc length) = ang1 maybe good to remember

OpenStudy (anonymous):

Good work, Amistre. You have calculated the intercepted arc, as they call it. The angle is half the intercepted arc.

OpenStudy (anonymous):

Do you understand terricherri.

OpenStudy (anonymous):

so they would be be 55 degrees right

OpenStudy (anonymous):

Yes, 'it' would be, the angle would be 55.

OpenStudy (anonymous):

but im stuck on this one same thing measure angle 1

OpenStudy (anonymous):

You have to recognize that straight line is the diameter. The diameter is the same degree as the arc. And follow from last problem.

OpenStudy (anonymous):

but how do i know will i know the arc of the circle if the degree is 1

OpenStudy (anonymous):

The degree is not 1, like you told me, it is just the name of the angle, like Angle A.

OpenStudy (anonymous):

nooooo what i mean was how will i know the arc if the degree in the diagram is 1

OpenStudy (anonymous):

The arc is equal to angle at the center of the circle. The angle is a straight line. It is kind of a trick question, yet very simple. What is the degree of a straight line?

OpenStudy (anonymous):

180 degrees

OpenStudy (anonymous):

Good. So the same way you determined the last one earlier. The arc was 110 and you determined the angle is 55. So the arc is 180, therefore the angle is...

OpenStudy (anonymous):

so it would be 110 - 55 -180?

OpenStudy (anonymous):

No, 110 and 55 is from the previous problem. It has nothing to do with this problem. It is just that you have to remember what you did in the last problem in order to do the others. The arc is 180, so the angle is half of 180.

OpenStudy (anonymous):

90

OpenStudy (anonymous):

Good

OpenStudy (anonymous):

Hmmmm

OpenStudy (anonymous):

what ?

OpenStudy (anonymous):

I asked if you wanted to meet me in my classroom for another session.

OpenStudy (anonymous):

You never replied...

OpenStudy (anonymous):

oh sorry i didnt know u were talking to me like directly cus they were aot of ppl in the group chat so ididnt know

OpenStudy (anonymous):

Yes, I was talking to you...

OpenStudy (anonymous):

0h ok